Can you teach me how to solve this problem.
The sum of three numbers is 105. The third is 11 less than ten times the second. Twice the first is 7 more than three times the second. Find the numbers.

Answer by stanbon(48568)

The sum of three numbers is 105. The third is 11 less than ten times the second. Twice the first is 7 more than three times the second. Find the numbers.
-----
Notice that the 1st and the 3rd are described in terms of the 2nd.
So. let the 2nd number be "x".
Then the 3rd is "10x-11"
Since 2(1st)=3x+7, 1st is (1/2)(3x+7)
EQUATION:
1st + 2nd + 3rd = 105
(1/2)(3x+7)+x+10x-11=105
(3/2)x+(7/2)+11x-11=105
(25/2)x=112.50
25x=225
x=9 (this is the 2nd number)
3rd number is 10x-11=79
1st number is (1/2)(3*9+7)=17
Cheers,
stan H.